OpenStudy (anonymous):

PLEASE HELP 48.) 1/3y+4=y-2 49.) 1/2(x-8)=5 50.) 1/7(p-35)=4 51.)3/4(b+4)=2/3(b-4)

thomaster (thomaster):

\(\large\dfrac{1}3y+4=y-2\) To get rid of the 1/3, multiply everything with 3

OpenStudy (anonymous):

3/3y+4(3)=y-2(3)

thomaster (thomaster):

You forgot the y on the right side

OpenStudy (anonymous):

3/3y+12=3y-6

thomaster (thomaster):

yes, 3/3 is the same as 1. So it'll be \(\large y+12=3y-6\)

thomaster (thomaster):

now bring y to the left and the numbers to the right

OpenStudy (anonymous):

1y+12=3y-6 1y-1y+12=3y-1y-6 12=2y-6 12-6+6=2y 12=2y y=6???

thomaster (thomaster):

This is wrong: 12=2y-6 12-6+6=2y 12=2y It should be 12=2y-6 12+6=2y 18=2y

OpenStudy (anonymous):

answer y=9??

thomaster (thomaster):

Yep :)

OpenStudy (anonymous):

could u just give me answers for 49-50 then you could help with 51-56???

OpenStudy (anonymous):

im exhausted i have to get up at 6am and i still have to get a shower ):

thomaster (thomaster):

Sorry but that's not how it works :( 1/2(x-8)=5 is very easy just distribute the 1/2 1/2x-4=5 Then multiply everything with 2 to get rid of the fraction x-8=10 x=18 Now do the same thing for 1/7(p-35)=4

OpenStudy (anonymous):

1/7p-7(35)=4(7) 7p-245=28 ????? thanks so much!

thomaster (thomaster):

Not entirely correct :( 1/7(p-35)=4 1/7p-5=4

OpenStudy (anonymous):

how do u make the 1/7 a whole number??

thomaster (thomaster):

multiply everything with 7

OpenStudy (anonymous):

so 7p-7(5)=4(7)

thomaster (thomaster):

no just p, not 7p

OpenStudy (anonymous):

7p-35=28

OpenStudy (anonymous):

p-35=28

OpenStudy (anonymous):

p-35=28+35 p= 63??

thomaster (thomaster):

correct :)

OpenStudy (anonymous):

your helping me learn (: p.s. i am exhausted tho lol

thomaster (thomaster):

I'm tired, it's 3:31 am here :P

OpenStudy (anonymous):

51.) 3/4(b+4)=2/3(b-4) JESUS!!!!!! its 9:32 pm here

thomaster (thomaster):

Oke last one: \(\large\dfrac{3}4(b+4)=\dfrac{2}3(b-4)\) First distribute 3/4 and 2/3

thomaster (thomaster):

This can be a little tricky. \(\large\dfrac{3}4*4=\dfrac{3}4*\dfrac{4}1-\dfrac{3*4}{4*1}=\dfrac{12}4\)

OpenStudy (anonymous):

12/4=3/1=3 answer=3?

thomaster (thomaster):

So when you distribute the left side of the equation, you'll have \(\large\dfrac{3}4b+3\)

thomaster (thomaster):

now the right side

OpenStudy (anonymous):

jeez im a lil lost 3/4 is confusing

thomaster (thomaster):

I know it's confusing but it's actually really simple. When you want to multiply fractions, always: \(\large\dfrac{a}b*\dfrac{c}d=\dfrac{a*c}{b*d}\) you can write any number (let's say 7) as \(\large\dfrac{7}1\) so \(\large\dfrac{3}4*4=\dfrac{3}4*\dfrac{4}1=\dfrac{3*4}{4*1}=\dfrac{12}4=3\)

OpenStudy (anonymous):

the left side is 3/4b+4 in the problem

OpenStudy (anonymous):

answer 3?

thomaster (thomaster):

left side = 3/4b+3

thomaster (thomaster):

the right side \(\large\dfrac{3}2*-4=\dfrac{3}2*\dfrac{-4}1-\dfrac{3*-4}{2*1}=\dfrac{-12}2\)

OpenStudy (anonymous):

-6 answer?

OpenStudy (anonymous):

i understand your version

thomaster (thomaster):

Yes :) So the new equation is: \(\large\dfrac{3}4b+3=\dfrac{2}3b-6\)

OpenStudy (anonymous):

thanks 5 last questions before you turn in??

thomaster (thomaster):

Now multiply the denominators (4*3=12) So first \(\large\dfrac{3}4b*12=\dfrac{3}4b*\dfrac{12}1\)

OpenStudy (anonymous):

36b/4 ?

thomaster (thomaster):

you can leave the b, he'll come back later. 36/4 = 9 So 9b

OpenStudy (anonymous):

9b=-6

thomaster (thomaster):

Now \(\large\dfrac{2}3b*12=\dfrac{2}3b*\dfrac{12}1\)

OpenStudy (anonymous):

24b/3

thomaster (thomaster):

24/3=8b don't forget to multiply +3 and -6 with 12 So the new equation is: \(\large 9b+36=8b-72\)

thomaster (thomaster):

Bring b to the left, numbers to the right and we have the answer :)

OpenStudy (anonymous):

17b+36

thomaster (thomaster):

No you have to subtract 8b

OpenStudy (anonymous):

1b=36 b=36 answer=b=36??

thomaster (thomaster):

Also subtract 36 So it'll be b=-108 Normally you would have to check if the answer is right with big solutions like these but I don't have time for that now

OpenStudy (anonymous):

52.)2d-6=1/4d+1

thomaster (thomaster):

I have to go now, post that one in a new question and i'm sure somebody will help you.

OpenStudy (anonymous):

ok thanks for all your help. VERY much appreciated (:

OpenStudy (anonymous):

sweet dreams